OpenStudy (anonymous):

Can anyone please help me with this one? I am not sure how to do it? Think about a kayaking race, a round trip airplane flight, a shopper on an escalator, or a swimming competition. Choose an activity where a rational expression can represent the total or difference of time. For example, what is the time to go round trip on a hot air balloon ride? Use full sentences to describe the real world situation. Write the addition or subtraction rational expression to represent your situation. Simplify the expression, showing all the work or explaining the steps using complete sentences.

OpenStudy (anonymous):

So I assume that by difference in time that the problem is asking for you to formulate an equation that will ultimately give you time (t) when you input all of the variables. For instance the hot air balloon ride I believe should be a rate=distance/time equation. I don't know how in depth you are supposed to get but I'd imagine that is what they are looking for.

OpenStudy (anonymous):

I am not sure either but the entire question is above.

OpenStudy (anonymous):

Well I would go with Time (t)=Distance (d)/ Rate (r). This function would allow you to input the rate of speed of the person, balloon, airplane, etc... giving you the change in time

OpenStudy (anonymous):

Ok have to see if I remember how to do that

OpenStudy (anonymous):

Depending on how in depth the class is, all of the variables can be functions of time. What class is the for?

OpenStudy (anonymous):

Algebra and the teacher is pretty relaxed and cool, I think as long as I answer the following she will be good. Use full sentences to describe the real world situation. Write the addition or subtraction rational expression to represent your situation. Simplify the expression, showing all the work or explaining the steps using complete sentences.

OpenStudy (anonymous):

oh ok then for a round trip airplane flight it would be \[\Delta t=(time \to city 1)\times(2)\]

OpenStudy (anonymous):

\[\Delta t\] is simply the change in time

OpenStudy (anonymous):

can you please help me work out this problem with describing the real world situation and explainging the steps, using actual numbers?

OpenStudy (anonymous):

So the time it takes to make a round trip air plane flight is the time it takes for the airplane to get to its destination plus the time it takes to return from its destination back home. So in the case of a plane flying from Philadelphia to New York, A plane takes 45 minutes to fly to New York, therefore it also takes 45 minutes to return to Philadelphia. Thus the change in time is 45 + 45 = 90 minutes or 1.5 hours

OpenStudy (anonymous):

Ok I understand that.

OpenStudy (anonymous):

do you need to explain all of the situations or just one?

OpenStudy (anonymous):

It says to write the addition or subtraction rational expression to represent your situation. Simplify the expression, showing all the work or explaining the steps using complete sentences.

OpenStudy (anonymous):

ok so the way we have it, we aren't taking into account: taxing, refilling the gas tanks of the plane or any of those types of real world scenarios. We have used the simplest situation so it is already in a simple rational form as 90=45 + 45

OpenStudy (anonymous):

Ok but do we need to change it to take into consideration the taxing or refiling of the gas tanks of the plane?

OpenStudy (anonymous):

yes we do. the total time would then not be 90 we would just list it as t for now. then we could just make estimations as to how long each of those situations would take

OpenStudy (anonymous):

so what do you want time (t) to be? the total time you are actually flying between cities, or total time, period?

OpenStudy (anonymous):

should we add another 30 minutes or another hour for the taxing and refilling of the plane?

OpenStudy (anonymous):

probably an hour

OpenStudy (anonymous):

ok so another hour, which would make the time 2.5 hours instead of 1.5 hours.

OpenStudy (anonymous):

yes that is right

OpenStudy (anonymous):

So would we do it as 45 + 45 + 60 = 150 minutes or 2.5 hours.

OpenStudy (anonymous):

so time= time it takes getting to NY+ time taxing/refilling/etc. + time to get back to Philly which is time=45+60+45=150 minutes=2.5 hours

OpenStudy (anonymous):

Ok thank you so very very much for all of your help. I understand it better now. You are a life saver.

OpenStudy (anonymous):

you're welcome have a good day

OpenStudy (anonymous):

You too thanks.

OpenStudy (anonymous):

I dunno